OpenStudy (anonymous):

I GIVE MEDALS Find all the zeros of each polynomial function. f(x) = x4 – 5x2 – 36 a. ±3, ±2i b. ±3 only c. ±2i only d. none of the above

OpenStudy (anonymous):

@dhirensharma

OpenStudy (cwrw238):

Hint :- let y = x^2 then solve the quadratic for y

OpenStudy (eliasperez191):

it is y=x^2

OpenStudy (anonymous):

zeros = 9,-9,-2

OpenStudy (mathmate):

@ninaesb ??

OpenStudy (anonymous):

im sorry , that was wrong . it was from a different problem .

OpenStudy (cwrw238):

y^2 - 5y - 36 = 0 (y - 9)(y + 5) = 0 so solution set for y is {-5 ,9} so we have x^2 = 9 or x^2 = -5

OpenStudy (mathmate):

@cwrw238 was there a typo?

OpenStudy (cwrw238):

ooops s sorry its (x - 9)(x + 4) = 0 yea y = 9 or -4

OpenStudy (cwrw238):

so x = +/- sqrt 9 , +/- sqrt(-4)

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

i just got back

OpenStudy (cwrw238):

u sure? whats sqrt 9?

OpenStudy (anonymous):

oh 3

OpenStudy (cwrw238):

+/- 3

OpenStudy (anonymous):

and 2

OpenStudy (cwrw238):

what about +/- sqrt (-4) have you came across the value i ?

OpenStudy (cwrw238):

i = sqrt(-1)

OpenStudy (cwrw238):

it is not real but an imaginary number

OpenStudy (cwrw238):

so i^2 = -1 and 2i * 2i = 4i^2 = -4 and -2i * -2i = 4i^2 = -4

OpenStudy (anonymous):

and the square of that is 2

OpenStudy (cwrw238):

the square of what is 2?

OpenStudy (anonymous):

of 4 right?

OpenStudy (anonymous):

so it would be a

OpenStudy (cwrw238):

no the square of 2i is -4 and also th square of -2i is -4 so the square roots of -4 are 2i and -2i

OpenStudy (cwrw238):

so which option is the correct answer?

OpenStudy (anonymous):

oh so c

OpenStudy (cwrw238):

dont forget the other 2 zeros....

OpenStudy (mathmate):

\((2i)^2=(2^2 i^2)=4(-1)=-4\) \((-2i)^2=((-2)^2 i^2)=4(-1)=-4\)

OpenStudy (anonymous):

ok if its 4 then its none of the above

OpenStudy (cwrw238):

patience is a virtue...

OpenStudy (cwrw238):

its a chantz417

OpenStudy (anonymous):

i said that along time ago!!!

OpenStudy (cwrw238):

so you did...

OpenStudy (anonymous):

you told me no

OpenStudy (cwrw238):

the zeros are +/- 3 and +/- 2i

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (cwrw238):

sorry - i got confused when you talked about the square of 2 and of 4